OpenStudy (anonymous):

What is the equation of the line described below written in slope-intercept form? the line passing through point (0, 0) and parallel to the line whose equation is 3x + 2y - 6 = 0

OpenStudy (anonymous):

parallel lines have the same slope

OpenStudy (anonymous):

2y+3x = 0

OpenStudy (anonymous):

so wouldn't that make it y=2/3x?

OpenStudy (anonymous):

no but y = -2/3x

OpenStudy (anonymous):

darn! ok, that makes sense now though, thank you